# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v4

@@ 3 Bias Correction and Asymptotic Theory
4 Simulation
We conduct the simulation study in this section. We first illustrate that our procedure is capable of reducing the bias of the MLE as well as delivering likelihood-based tests with an empirical size very close to the nominal level. Followed by this, we then consider a different design in which the MLEs and the tests suffer from severe biases and size-distortions. We demonstrate that our approach leads to dramatic improvements on the estimators and substantial reductions of the Type-I error risks of the tests, even in a severely biased setting. Finally in this section, we compare our method with the Jackknife [^7] Since there has currently been no Jackknife procedures for DN-HP models. We compared our method to the Jackknife of Fernández-Val and Weidner (2016), who consider models with fixed effects only in the intercept. and show that our method out-performs.
In Design 1, we consider a dynamic binary response model as:
$Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ (4.1)
where the true value is $\theta_{0}=(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ and the disturbance $\varepsilon_{it}$ is standard logistic, independent and identically distributed (i.i.d.) over $(i,t)$ and independent from $X_{it}$. The true incidental parameter $\{\alpha_{1,i}^{0},\gamma_{1,t}^{0}\}$ are generated from $\mathcal{N}(0,0.01)$, and demeaned after they are generated. $\{\alpha_{2,i}^{0},\gamma_{2,t}^{0}\}$ are also generated from $\mathcal{N}(0,0.01)$ but only $\gamma_{2,t}^{0}$ is demeaned. Conditional on the generated $\alpha_{k,i}^{0}$ and $\gamma_{k,t}^{0}$ ($k=1,2$), the regressor $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$ i.i.d. over $(i,t)$ and the initial $Y_{i0}=\mathds{1}\{(\beta_{0}+\alpha_{1,i}^{0})X_{i0}+\alpha_{2,i}^{0}+\varepsilon_{i0}>0\}$ with $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$ i.i.d. over $(i,t)$. We also simulate a static model, where we set $\rho\equiv 0$ in (4.1), using data generated in the same way but setting $\rho_{0}=0$ (hence $\theta_{0}=\beta_{0}$).
Table 1 reports the estimation bias and RMSE, under Design 1, of $\widehat{\theta}_{L}$ and $\widehat{\theta}$ for both the static ($\theta_{0}=\beta_{0}$) and dynamic ($\theta_{0}=(\rho_{0},\beta_{0})^{\prime}$) logit models. For the static model, $\widehat{\beta}$ denotes the uncorrected MLE obtained by maximing $\widehat{l}(\theta)$, and $\widehat{\beta}_{L}$ is the corrected estimator, obtained by maximing $L(\theta)$. For the dynamic model, $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimator for $(\rho_{0},\beta_{0})^{\prime}$, while $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents the bias-corrected estimator, with the truncation parameter $\tau\in(1,2)$. For the static model (with $\rho_{0}\equiv 0$), we set $\tau=0$ since the data are independent over $i$, conditional on $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}$. The results reveal that the uncorrected estimators are severely biased in both models, while our bias-corrected estimators substantially reduce the bias. For instance, in the static model with $(N,T)=(30,30)$, the bias for $\widehat{\beta}_{L}$ is only $0.05$, which is less than $1/3$ of the significant bias of $0.158$ for $\widehat{\beta}$. As the sample size increases to $(N,T)=(90,90)$, $\widehat{\beta}_{L}$ is nearly unbiased, while the bias of $\widehat{\beta}$ remaining five times larger than that of $\widehat{\beta}_{L}$. Furthermore, the RMSEs from our corrected estimators are consistently lower than those of the uncorrected estimators, especially with smaller sample sizes. This indicates that our procedure is capable of reducing the IPP bias without increasing the RMSE. For large $(N,T)$, the choice of $\tau$ has negligible impact on the performance of our bias-corrected estimators.
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.158$ $0.253$ $0.075$ $0.109$ $0.051$ $0.072$
$\widehat{\beta}_{L}$ $0.050$ $0.175$ $0.018$ $0.073$ $0.010$ $0.049$
Dynamic Model
$\widehat{\beta}$ $0.165$ $0.261$ $0.074$ $0.111$ $0.046$ $0.071$
$\widehat{\beta}_{L}^{(1)}$ $0.059$ $0.181$ $0.017$ $0.077$ $0.006$ $0.051$
$\widehat{\beta}_{L}^{(2)}$ $0.061$ $0.183$ $0.018$ $0.078$ $0.007$ $0.051$
$\widehat{\rho}$ $-0.107$ $0.201$ $-0.048$ $0.093$ $-0.033$ $0.058$
$\widehat{\rho}_{L}^{(1)}$ $-0.044$ $0.159$ $-0.014$ $0.076$ $-0.010$ $0.047$
$\widehat{\rho}_{L}^{(2)}$ $-0.038$ $0.161$ $-0.009$ $0.077$ $-0.006$ $0.046$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$. For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$. The number of replications is $1000$.
Next, under Design 1, we conduct LR tests using both the corrected and the uncorrected likelihoods for the null hypothesis $H_{0}:\theta_{0}=(0.5,0.5)^{\prime}+\delta$ with $\delta\in\{0,\pm 0.2,\pm 0.15,\pm 0.1\}$. We report the empirical size ($\delta=0$) and the empirical power ($\delta\neq 0$) at the $5\%$ significance level. For the corrected likelihood, we set the truncation parameter $\tau=0$ for the static model and $\tau=1$ for the dynamic model. Table 2 presents the empirical size. The results show that the LR test using the uncorrected likelihood suffers from a severe size distortion. On the opposite, the empirical size from our bias-corrected likelihood maintains the significance level as the sample size increases.
Table 2: Empirical Size of LR Tests Based on Corrected and Uncorrected Likelihoods for Logit Model, Design 1
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $16\%$ $9\%$ $18\%$ $10\%$
$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level of $5\%$. The number of replications is $1000$.
Table 3 presents the empirical power of LR tests based on our corrected likelihood under Design 1. The empirical power increases as both the value of $\delta$ and the sample size grow. For $(N,T)=(90,90)$, when $\delta=\pm 0.2$, the rejection rates reach nearly $100\%$ for both the static and dynamic models. We also perform LM and Wald test, reporting the empirical size and power under the same setting in Appendix B.1. The results align with those of the LR test. We additionally perform the simulations for a probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$ i.i.d. over $(i,t)$ and independent from $X_{it}$. The results are provided in Appendix B.2.
Table 3: Empirical Power of LR Test for Logit Model using $L(\theta)$, Design 1
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $45\%$ $32\%$ $23\%$ $10\%$ $14\%$ $21\%$
$(60,60)$ $86\%$ $63\%$ $37\%$ $25\%$ $49\%$ $74\%$
$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
Dynamic Model
$(30,30)$ $48\%$ $33\%$ $21\%$ $18\%$ $28\%$ $43\%$
$(60,60)$ $95\%$ $74\%$ $40\%$ $45\%$ $75\%$ $94\%$
$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
Next, we conduct the same experiment under Design 2, where we generate $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$, keeping everything else intact as in (4.1). We increase the volatility of $\alpha_{k,i}^{0}$, $\gamma_{k,t}^{0}$, $X_{it}$, and $X_{i0}$ to induce larger biases and size distortions. This time, we present the results for the probit model and leave that for the logit model in Appendix B.1.
Table 4 presents the comaprison of the bias and RMSE of the corrected and the uncorrected estimators for the probit model, under Design 2. First of all, the probit estimator here is much more biased than the previous logit counterparts. For the static model with $(N,T)=(30,30)$, $\widehat{\beta}$ contains a bias of $0.943$. Recall that the true value is only $\beta_{0}=0.5$, implying that the bias is almost $200\%$. Our corrected estimator $\widehat{\beta}_{L}$ successfully reduces this bias to $0.328$, which is one third of the uncorrected and roughly $60\%$ biased. As the sample sizes increase, the bias systematically becomes smaller both for the uncorrected and corrected estimators, with dramatic improvements observed from the corrected ones both for the stiatic and dynamic models. We indeed observe that the RMSE of $\widehat{\beta}_{L}$ is twice as large as that of $\widehat{\beta}$ for the static model under $N,T=(30,30)$. However, it is the only incidence here and we attribute this to some sort of numerical coincidence [^8] Below, when we compare our method with the Jackknife under the same setting, it will be seen that the RMSE of our corrected estimator is indeed smaller than the uncorrected. See also Footnote 9 below.. It can be checked that the RMSE of our corrected estimator is no worse the uncorrected one in every other case.
Table 4: Comparisons of Uncorrected and Corrected Probit Estimates, Design 2
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.943$ $2.272$ $0.151$ $0.252$ $0.073$ $0.081$
$\widehat{\beta}_{L}$ $0.328$ $5.170$ $0.041$ $0.062$ $0.017$ $0.040$
Dynamic Model
$\widehat{\beta}$ $0.909$ $1.842$ $0.166$ $0.248$ $0.084$ $0.100$
$\widehat{\beta}_{L}^{(1)}$ $0.207$ $0.396$ $0.048$ $0.078$ $0.022$ $0.042$
$\widehat{\beta}_{L}^{(2)}$ $0.202$ $0.282$ $0.049$ $0.080$ $0.022$ $0.042$
$\widehat{\rho}$ $-0.030$ $0.162$ $-0.016$ $0.073$ $-0.011$ $0.045$
$\widehat{\rho}_{L}^{(1)}$ $-0.015$ $0.140$ $-0.005$ $0.066$ $-0.004$ $0.044$
$\widehat{\rho}_{L}^{(2)}$ $-0.009$ $0.141$ $-0.001$ $0.067$ $-0.001$ $0.045$
[…]
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$. For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$. The number of replications is $1000$.
Table 5 below reports the empirical size of the LR test for the probit model under Design 2. As before, the null hypothesis is $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model and $H_{0}:\beta_{0}=0.5$ for the static. We set the truncation parameter $\tau=0$ for the static model and $\tau=1$ for the dynamic model. Here the size distortion is much more fatal than in the logit case above. For the static model with $(N,T)=(30,30)$, the uncorrected LR test produces an empirical size of $77\%$. Even under $(N,T)=(90,90)$, the empirical size is still $72\%$, indicating a prohibitively high risk of committing a Type-I error. On the other hand, the $L(\theta)$-based LR test delivers a $14\%$ empirical size only, which is almost a $60\%$ decrease in magnitude. Similar performance is observed in other cases. The results here demonstrate that our method can reduce the type-I error risks by a major margin. Indeed, our test still over-rejects under $(N,T)=(90,90)$. This is anticipated, because the IPP under Design 2 is much more drastic and $(N,T)=(90,90)$ may not be sufficient for the test statistic to be well-approximated by $\chi^{2}$. However, it is clear that, as $(N,T)$ increase, the empirical size does become smaller. We have tried additional small-scale simulations (not reported though) with even larger $(N,T)$ and found that the empirical size still decreases with $(N,T)$. We similarly report the results from the LM and the Wald test in Appendix B.2. We report the empirical power under Design 2 in the Appendix B, since they do not add too much value in the discussion here. A quick summary for those results is that our procedure produces tests with stronger power in Design 2.
Table 5: Empirical Size of LR Tests Based on Corrected and Uncorrected Likelihoods for Logit Model, Design 2
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $77\%$ $39\%$ $71\%$ $37\%$
$(60,60)$ $76\%$ $20\%$ $66\%$ $19\%$
$(90,90)$ $72\%$ $14\%$ $63\%$ $14\%$
[…]
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level of $5\%$. The number of replications is $1000$.
Finally, we compare our method with the Jackknife of Fernández-Val and Weidner (2016). We consider the same probit model and continue the same setting and design as above. Table 6 displays the estimation bias and the RMSE of the uncorrected estimators $\{\widehat{\rho},\widehat{\beta}\}$, our corrected estimator $\{\widehat{\rho}_{L},\widehat{\beta}_{L}\}$, and the Jackknife-corrected estimator $\{\widehat{\rho}_{J},\widehat{\beta}_{J}\}$ for the dynamic model. For the static model, only $\widehat{\beta}$, $\widehat{\beta}_{L}$, and $\widehat{\beta}_{J}$ apply and are reported. We set the truncation parameter $\tau=0$ for static model and $\tau=1$ for dynamic model. Across all pairs of $(N,T)$, the Jackknife-corrected $\widehat{\beta}_{J}$ consistently exhibits larger bias and RMSE than our $\widehat{\beta}_{L}$. Even worse, $\widehat{\beta}_{J}$ shows a larger bias and RMSE than the uncorrected $\widehat{\beta}$ across all pairs of $(N,T)$. For instance, with $(N,T)=(30,30)$ and the static model, $\widehat{\beta}_{J}$ shows a bias of $-4.929$, which is roughly $900\%$ biased (in magnitude) relative to the true value $\beta_{0}=0.5$. Observe that the uncorrected estimator $\widehat{\beta}$ only shows a roughly $170\%$ bias relative to the true value. Our corrected $\widehat{\beta}_{L}$ performs the best among the three, showing a $36\%$ bias [^9] Also notice that the RMSE of $\widehat{\beta}_{L}$ in Table 6 is much smaller than that of $\widehat{\beta}_{L}$ in Table 4, where both tables are essentially under the same setting. This motivates our attribution of the large RMSE in Table 4 to numerical coincidence.. To summarize, we find that the Jackknife is not a suitable candidate for DN-HP models. Although it does reduce the bias for the homogeneous coefficient for large $N,T$, the bias the Jackknife causes to the heterogeneous parameter is dangerously large, especially in small sample.
Table 6: Comparison with the Jackknife Bias Correction Method
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.846$ $2.168$ $0.151$ $0.261$ $0.074$ $0.082$
$\widehat{\beta}_{L}$ $0.182$ $0.339$ $0.041$ $0.068$ $0.019$ $0.033$
$\widehat{\beta}_{J}$ $-4.929$ $14.383$ $-0.504$ $0.981$ $-0.105$ $0.214$
Dynamic Model
$\widehat{\beta}$ $1.040$ $2.343$ $0.184$ $0.287$ $0.080$ $0.090$
$\widehat{\beta}_{L}$ $0.209$ $0.332$ $0.052$ $0.082$ $0.020$ $0.049$
$\widehat{\beta}_{J}$ $-3.352$ $7.708$ $-0.558$ $1.125$ $-0.118$ $0.217$
$\widehat{\rho}$ $-0.046$ $0.169$ $-0.022$ $0.072$ $-0.014$ $0.046$
$\widehat{\rho}_{L}$ $-0.028$ $0.145$ $-0.011$ $0.067$ $-0.009$ $0.044$
$\widehat{\rho}_{J}$ $-0.029$ $0.169$ $-0.006$ $0.068$ $-0.002$ $0.044$
[…]

### v5

@@ C.3 Asymptotic Theory
D Relation to Existing Approaches
As discussed above, our method is a non-trivial extension of Bester and Hansen (2009) and Arellano and Hahn (2016) to two-way DN-HP models. In this section, we explain the technical challenges in developing a likelihood-based bias correction approach for the DN-HP models and provide Example D.1 to illustrate how our method relates to the aforecited. We also leave Remark 7 to briefly discuss the difference of our procedure to some existing methods providing fixed-$T$ consistency for the logit model
The primary technical challenge lies in the high dimensionality of the two-way DN-HP models, due to incidental parameter vector $\phi$ (or $\psi$) having a dimension increasing with $N$ and $T$. In what follows, we briefly re-visit the procedure of Arellano and Hahn (2016), pointing out the key difference, and explain the challenges. For simplicity, we assume that the fixed effects are only scalars [^8] We will briefly discuss multi-dimensional fixed effects later.. Consider an arbitrary panel data model with individual effects only, having likelihood function
$\ell\left(\theta,\alpha\right):=\frac{1}{N}\sum_{i=1}^{N}l_{i}\left(\theta,\alpha_{i}\right),\qquad l_{i}\left(\theta,\alpha_{i}\right):=\frac{1}{T}\sum_{i=1}^{T}\log f\left(Y_{it}|X_{it};\theta,\alpha_{i}\right),$
where $f\left(Y_{it}|X_{it};\theta,\alpha_{i}\right)$ is a known density and $\alpha:=\left(\alpha_{1},\ldots,\alpha_{N}\right)^{\prime}$. It is important to note that $l_{i}\left(\theta,\alpha_{i}\right)$ are independent across $i$, in the sense that $l_{i}\left(\theta,\alpha_{i}\right)$ and $l_{j}\left(\theta,\alpha_{j}\right)$ share no common arguments except $\theta$. Because of such independence, an expansion similar to (2.5) can be established individual by individual, as
$\quad\quad l_{i}\left(\theta,\widehat{\alpha}_{i}\left(\theta\right)\right)$
$=l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)+\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)\left(\widehat{\alpha}_{i}\left(\theta\right)-\alpha_{i}\left(\theta\right)\right)+\frac{1}{2}\partial_{\alpha_{i}^{2}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)\left(\widehat{\alpha}_{i}\left(\theta\right)-\alpha_{i}\left(\theta\right)\right)^{2}+r_{i}^{l}\left(\theta\right),$ (D.1)
where, for some $c_{i}\in\left(0,1\right)$,
$r_{i}^{l}\left(\theta\right)$ $:=\frac{1}{6}\partial_{\alpha_{i}^{3}}l_{i}\left(\theta,c_{i}\alpha_{i}\left(\theta\right)+\left(1-c_{i}\right)\alpha_{i}\left(\theta\right)\right)\left(\widehat{\alpha}_{i}\left(\theta\right)-\alpha_{i}\left(\theta\right)\right)^{3},$
$\widehat{\alpha}_{i}\left(\theta\right)$ $:=\arg\max_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\right),\qquad\alpha_{i}\left(\theta\right):=\arg\max_{\alpha_{i}}\mathbb{E}l_{i}\left(\theta,\alpha_{i}\right),$ (D.2)
$\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)$ denotes $\left[\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\right)\right]|_{\alpha_{i}=\alpha_{i}\left(\theta\right)}$, and $\partial_{\alpha_{i}^{2}}l_{i}\left(\cdot,\cdot\right)$ and $\partial_{\alpha_{i}^{3}}l_{i}\left(\cdot,\cdot\right)$ are defined analogously. Here, notice importantly that $\widehat{\alpha}_{i}\left(\theta\right)$ and $\alpha_{i}\left(\theta\right)$ are obtained by solving, respectively, $\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\right)=0$ and $\mathbb{E}\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\right)=0$ individually. This is possible due to the independence of $l_{i}\left(\theta,\alpha_{i}\right)$ across $i$. Subsequently, a similar expansion to (2.6) can be constructed, again individually, as
$0$ $=\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)+\partial_{\alpha_{i}^{2}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)\left(\widehat{\alpha}_{i}\left(\theta\right)-\alpha_{i}\left(\theta\right)\right)+r_{i}^{s}\left(\theta\right)$ (D.3)
$\widehat{\alpha}_{i}\left(\theta\right)-\alpha_{i}\left(\theta\right)$ $=-\frac{1}{\partial_{\alpha_{i}^{2}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)}\left(\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)+r_{i}^{s}\left(\theta\right)\right),$
where, for some $d_{i}\in\left(0,1\right)$,
$r_{i}^{s}\left(\theta\right)=\frac{1}{2}\partial_{\alpha_{i}^{3}}l_{i}\left(\theta,d_{i}\alpha_{i}\left(\theta\right)+\left(1-d_{i}\right)\alpha_{i}\left(\theta\right)\right)\left(\widehat{\alpha}_{i}\left(\theta\right)-\alpha_{i}\left(\theta\right)\right)^{2}.$
Similarly combining (D.3) and (D.1), we have
$l_{i}\left(\theta,\widehat{\alpha}_{i}\left(\theta\right)\right)=l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)-\frac{1}{2}\frac{\left[\partial_{\alpha_{i}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)\right]^{2}}{\partial_{\alpha_{i}^{2}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)}-r_{i}\left(\theta\right)$
for a remainder term $r_{i}\left(\theta\right)$ depending on $r_{i}^{s}\left(\theta\right)$ and $r_{i}^{l}\left(\theta\right)$, amongst other things. The subsequent steps are omitted because they are similar to their DN-HP counterparts.
The steps above are all followed only individually. This is possible due to the independence of $l_{i}\left(\theta,\alpha_{i}\right)$; and the benefit is that $r_{i}\left(\theta\right)$, $\partial_{\alpha_{i}^{2}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)$, etc. are only scalars. For two-way DN-HP models, the likelihood $l_{i,t}(\theta,\phi)=\log f(Y_{it}|X_{it},\theta+\alpha_{i}+\gamma_{t})$ (defined in 2.2) depends on both $\alpha_{i}$ and $\gamma_{t}$. This is unlike models with individual effects only, and it implies that $\widehat{\psi}\left(\theta\right)$ must be obtained by solving the system of equations $\partial_{\psi}l\left(\theta,\psi\right)=0$ jointly, unlike $\widehat{\alpha}_{i}\left(\theta\right)$. It further implies that the expansions in (2.5) and (2.6) must be carried out jointly w.r.t. the entire vector $\psi$. In other words, these expansions are high-dimenional for DN-HP models. Consequently, it is much more involved to develop a likelihood-based bias correction for two-way DN-HP models, in the follows notable ways, all due to the high-dimensionality.
  1. 1.
    For models with individuals effects only, $r_{i}^{s}\left(\theta\right)$ is only a scalar and it is relatively straightforward to establish its rate of convergence. On the other hand, its DN-HP counterpart, given in (E.6), is a high-dimensional vector, with each component itself being a quadratic form of $\widehat{\psi}\left(\theta\right)-\psi(\theta)$ and $G_{\psi_{j}}(\theta,a_{j})$ that are already high-dimensional objects. It is technically more challenging to control the rate of convergence of (E.6). For instance, one obvious issue is that the choice of norm becomes relevant for (E.6) but not $r_{i}^{s}\left(\theta\right)$. The same arguments apply to other remainder terms.
  2. 2.
    When the model only contains (scalar) individual effects, the incidental parameter Hessian $\partial_{\alpha_{i}^{2}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)$ is only a scalar and the rate of convergence of its inverse, $1/\partial_{\alpha_{i}^{2}}l_{i}\left(\theta,\alpha_{i}\left(\theta\right)\right)$, is relatively straightforward. For the DN-HP models, however, the incidental parameter Hessian $H\left(\theta\right)$ is a high-dimensional matrix. It is more complex to control the rate of its inverse, $\left[H\left(\theta\right)\right]^{-1}$.
  3. 3.
    When only individual effects are present, there is only one scalar bias term. It is relatively straightforward to derive its rate of convergence and propose an estimator for it. On the contrary, for the DN-HP models, the bias correction term $B(\theta)$, defined in (2.7), is a quadratic form of $s\left(\theta\right)$ and $H\left(\theta\right)$ which are high-dimensional objects. In addition to this, we show that $B(\theta)\$can be approximated by $B_{\alpha}(\theta)+B_{\gamma}(\theta)$, which is not needed for models with individual effects only.
Other challenges include, e.g., the lack of stationarity across $t$ (due to the existence of $\gamma_{t}^{0}$), the need for the normalization and reparameterization, etc. Note that, for models with individual effects only, even if $\alpha_{i}$ is multi-dimensional, the steps above can still be carried out individually, implying that the above-discussed high-dimensionality problems still would not happen, as long as the dimension of $\alpha_{i}$ is fixed.
Next, in Example D.1 below, we illustrate how our approach reduces to Arellano and Hahn (2016) and apply it to a dynamic logit model with only scalar individual effects in the intercept.
Example D.1 (dynamic logit with individual effects only)
Consider a model with only scalar individual effects in the intercept, whose likelihood for individual $i$ at time $t$ is $l_{i,t}\left(\theta,\alpha_{i}\right)$, and assume that the data is independent over $i$ while stationary across $t$. Since, reparameterization is not needed, we work with the original likelihood $\ell\left(\theta,\phi\right)\equiv N^{-1}T^{-1}\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}\left(\theta,\alpha_{i}\right)$ for $\phi\equiv\left(\alpha_{1},\ldots,\alpha_{N}\right)^{\prime}$. Given that $l_{i,t}\left(\theta,\alpha_{i}\right)$ are independent across $i$, $\widehat{\phi}(\theta)$ and $\phi(\theta)$ can be constructed individually, as
$\widehat{\phi}(\theta)\equiv\left(\widehat{\alpha}_{1}\left(\theta\right),\ldots,\widehat{\alpha}_{N}\left(\theta\right)\right)^{\prime},\qquad\phi(\theta)\equiv\left(\alpha_{1}\left(\theta\right),\ldots,\alpha_{N}\left(\theta\right)\right)^{\prime},$
where, $\widehat{\alpha}_{i}\left(\theta\right)$ and $\alpha_{i}\left(\theta\right)$ are defined in (D.2). Next, notice that $\mathbb{E}s_{i,t}^{\alpha}(\theta,\alpha_{i}(\theta))=\mathbb{E}T^{-1}\sum_{t=1}^{T}s_{i,t}^{\alpha}(\theta,\alpha_{i}(\theta))=0$, where the first equality is because of stationarity and the second because $\alpha_{i}(\theta)$ maximizes $\mathbb{E}l_{i}\left(\theta,\alpha_{i}\right)$. This implies that $S_{\alpha\alpha}(\theta)$ is a diagonal matrix defined as
$S_{\alpha\alpha}(\theta)=\frac{1}{N^{2}T}\operatorname*{diag}\left\{S_{i}^{\alpha\alpha}\left(\theta\right)\right\},\qquad S_{i}^{\alpha\alpha}\left(\theta\right):=\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left\{s_{i,t}^{\alpha}(\theta,\alpha_{i}(\theta))s_{i,s}^{\alpha}(\theta,\alpha_{i}(\theta))\right\}.$
Similarly, the expected Hessian is also diagonal, as
$\mathbb{E}H(\theta)=\frac{1}{N}\operatorname*{diag}\left\{\mathbb{E}h_{i}^{\alpha\alpha}(\theta)\right\},\qquad h_{i}^{\alpha\alpha}(\theta):=\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{i}^{2}}l_{i,t}\left(\theta,\alpha_{i}\left(\theta\right)\right).$
By stationarity, the inverse of $\mathbb{E}H(\theta)$ is simply
$H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)\equiv[\mathbb{E}H(\theta)]^{-1}=N\operatorname*{diag}\left\{[\mathbb{E}\partial_{\alpha_{i}^{2}}l_{i,t}\left(\theta,\alpha_{i}\left(\theta\right)\right)]^{-1}\right\}.$
Finally, seeing $B_{\gamma}(\theta)\equiv 0$ without time effects and using the properties of $\operatorname*{trace}$, we have
[…]